# Jean DELHAYE

### v2

@@ 1. Introduction
Let $(X_{N})_{N\in\mathbb{N}}$ be a sequence of irreducible, aperiodic, finite-state Markov chains, with $\mu_{N}(t)$ denoting the distribution of $X_{N}$ after $t$ steps and $\mu_{N}(\infty)$ its stationary distribution. Define the total variation distance to equilibrium at time $t$ as
$d_{N}(t)=d_{\mathrm{TV}}(\mu_{N}(t),\mu_{N}(\infty)),$
where the total variation distance $d_{\mathrm{TV}}(\mu,\nu)$ between two probability measures $\mu$ and $\nu$ on a compact set $K$ is given by
$d_{\mathrm{TV}}(\mu,\nu)=\sup_{A\subset K}|\mu(A)-\nu(A)|,$
where the supremum is taken over all Borel sets. Consider a sequence of times $(t_{N})_{N\in\mathbb{N}}$. We say that $(X_{N})_{N\in\mathbb{N}}$ exhibits a cutoff in total variation distance at time $(t_{N})_{N\in\mathbb{N}}$ if, for every $\epsilon>0$,
$\lim_{N\to\infty}d_{N}((1-\epsilon)t_{N})=1\quad\text{and}\quad\lim_{N\to\infty}d_{N}((1+\epsilon)t_{N})=0.$
This indicates that the convergence to equilibrium undergoes a sharp transition, collapsing from $1$ to 00 around the time $t_{N}$.
@@
The first example of the cutoff phenomenon in groups comes from the work of P. Diaconis and M. Shahshahani in the 1980s on finite groups [15]. This groundbreaking discovery sparked significant interest in the phenomenon, leading to the identification of numerous examples across various contexts. Notable cases include Brownian motion on simple compact Lie groups and symmetric spaces [28], as well as random walks on random graphs [26]. In this paper, we focus on quantum compact groups, an object initially introduced by S.L. Woronowicz in [36]. The representation theory of compact groups extends naturally to the quantum setting, enabling the application of similar techniques. This has been studied extensively, starting with finite quantum groups, as explored by J.P. McCarthy in [27], and extended to infinite compact quantum groups in works such as [20, 21].
To gain deeper insight into this phenomenon, we zoom in on the interval where this sharp transition occurs. The cutoff phenomenon suggests that the width of this interval is negligible compared to the sequence $(t_{N})_{N\in\mathbb{N}}$, prompting us to investigate the next significant higher-order term. Specifically, if there exists a sequence $(s_{N})_{N\in\mathbb{N}}$ and a continuous function $f$ that decreases from $1$ to 00 such that, for all $c\in\mathbb{R}$,
$\lim_{N\to\infty}d_{N}(t_{N}+cs_{N})=f(c),$
then $f$ is referred to as the cutoff profile or limit profile of $(X_{N})_{N\in\mathbb{N}}$.
@@
While determining a cutoff profile is inherently challenging, there are several processes for which such profiles have been established. For instance, in the case of discrete processes like random walks on finite groups, the random transposition shuffle [33] exhibits the following asymptotic behavior:
@@
$d_{N}\left(\frac{N}{2}\ln N+cN\right)\underset{N\to\infty}{\longrightarrow}d_{\mathrm{TV}}\left(\mathrm{Pois}(1+e^{-c}),\mathrm{Pois}(1)\right),\quad c\in\mathbb{R}.$
@@
Other examples include the lazy random walk on the hypercube [31], the dovetail shuffle [5], the simple exclusion process on the circle [24], the Ehrenfest Urn model with multiple urns [30], the Gibbs Sampler [30], and random cycles [30].
The study of cutoff profiles on compact quantum groups is relatively recent and primarily based on the work [23], where the authors compute the profile for the Brownian motion on the quantum orthogonal group $O_{N}^{+}$ and the quantum permutation group $S_{N}^{+}$, as well as for quantum random transpositions on $S_{N}^{+}$.
@@
While the Brownian motion is the most natural continuous process on simple compact Lie groups, defining a clear analog on most compact quantum groups is not straightforward. The orthogonal and permutation quantum groups provide natural processes that can be interpreted as such (stemming from analogies between the central generating functionals on $O_{N}^{+}$ and $S_{N}^{+}$, as proved in [12, Thm 10.2] and [18, Thm 10.10], and the classical case for compact Lie groups in [25]). However, extending this concept to other compact quantum groups presents significant challenges.
@@
Our main result is the identification and analysis of the Brownian motion on the unitary quantum group $U_{N}^{+}$, which is detailed in Section 3, along with the partial computation of its limit profile, presented in Section 5. Specifically, we establish in Corollary 5.2 that, for appropriate extensions $\widetilde{d}_{N}$ of the distances $d_{N}$ in the quantum setting, we have
@@
$\widetilde{d}_{N}\left(N\ln(\sqrt{2}N)+cN\right)\underset{N\to\infty}{\longrightarrow}d_{\mathrm{TV}}\left(\mathrm{Meix}^{+}(-e^{-c},0)*\delta_{e^{-c}},\nu_{\mathrm{SC}}\right),\quad c\geq 0,$
@@
$\underset{N\to\infty}{\lim\sup}\,\widetilde{d}_{N}\left(N\ln(\sqrt{2}N)+cN\right)\geq d_{\mathrm{TV}}\left(\mathrm{Meix}^{+}(-e^{-c},0)*\delta_{e^{-c}},\nu_{\mathrm{SC}}\right),\quad c<0,$
@@
where $\nu_{\mathrm{SC}}$ denotes the semicircle distribution and $\mathrm{Meix}^{+}$ refers to the free Meixner distribution (see Subsection 2.2 for definitions).
@@
where $\nu_{c}$ represents a limit measure (or state) for the process at time $t_{N}+cs_{N}$. The main difficulty in the quantum setting arises because the process is not absolutely continuous with respect to the stationary distribution whenever $c<c_{0}$ (for some $c_{0}\in(0,1)$, depending on the scaling of the sequences). This phenomenon, first noted in [23], shows that processes exhibit an atom in certain regions. In our case, the central algebra of the quantum unitary group $U_{N}^{+}$ is not commutative, unlike $O_{N}^{+}$ and $S_{N}^{+}$. Additionally, in the region where absolute continuity is lost, the singular part does not reduce to a single atom, but rather involves a more complex structure that we were unable to fully identify, which is why we can only establish an lower bound in this region.
The article concludes in Section 6, where we discuss the general observation that cutoff profiles tend to be similar across different structures. This raises the question of whether the structure alone determines the cutoff profile. We address this by providing an example of a process on $O_{N}^{+}$ that exhibits a distinct limit profile from that of Brownian motion.
Acknowledgement
@@
For more information, we encourage the reader to consult [13] for an excellent source on mixing times, and [22], [29], or [4] for a detailed study of compact quantum groups and their representation theory.
@@ 2. Preliminaries
In this paper, we will focus on compact quantum groups, which may be unfamiliar to probabilist readers. To address this, we will dedicate a preliminary section to defining and presenting key aspects. Specifically, we will introduce free unitary and orthogonal quantum groups, along with relevant results regarding Lévy processes on these objects.
@@ 2.1. Free unitary quantum groups
Free unitary quantum groups, a specific class of compact quantum groups, were introduced by S. Wang in [35]. Quantum groups, more broadly, were originally defined by S.L. Woronowicz [36]. These structures involve C*-algebras, fitting their noncommutative topological nature. However, in this article, we will present an alternative definition that emphasizes the algebraic aspects, making it more accessible to non-expert readers.
Let us recall that a *-algebra is a complex unital algebra $A$ endowed with an involution $A\to A,x\mapsto x^{*}$, i.e. an antimultiplicative antilinear map such that $x^{**}=x$ for all $x\in A$. Also, a *-ideal $B$ of $A$ is a *-subalgebra such that $aB\cup Ba\subset B$ for all $a\in A$.
@@
In other words, $\mathcal{O}(U_{N}^{+})=\mathbb{C}\langle u_{ij}\rangle_{1\leq i,j\leq N}/I$ where $\mathbb{C}\langle u_{ij}\rangle_{1\leq i,j\leq N}$ denotes the *-algebra of noncommutative polynomials in variables $u_{ij}$, $u_{ij}^{*}$ ($1\leq i,j\leq N$) and $I$ denotes the *-ideal generated by the elements
$\left\{\sum_{k=1}^{N}u_{ik}u_{jk}^{*}-\delta_{ij},\sum_{k=1}^{N}u_{ki}^{*}u_{kj}-\delta_{ij},\sum_{k=1}^{N}u_{ik}^{*}u_{jk}-\delta_{ij},\sum_{k=1}^{N}u_{ki}u_{kj}^{*}-\delta_{ij}\right\}.$
Let $U_{N}$ be the usual unitary group, $c_{ij}:U_{N}\to\mathbb{C}$ the function sending a matrix to its $(i,j)$-th coefficient and let $\mathcal{O}(U_{N})$ be the algebra of regular functions on $U_{N}$, i.e. the *-algebra generated by the functions $c_{ij}$, where the involution corresponds to the complex conjugation: $c_{ij}^{*}=\overline{c}_{ij}$. Then, quotienting $\mathcal{O}(U_{N}^{+})$ by its commutator ideal yields a surjection
$\pi\left\{\begin{matrix}\mathcal{O}(U_{N}^{+})&\longrightarrow&\mathcal{O}(U_{N})\\ u_{ij}&\longmapsto&c_{ij}.\end{matrix}\right.$
In that sense, $U_{N}^{+}$ can be thought of as a noncommutative version of $U_{N}$. The group structure of $U_{N}$ can be recovered in an algebraic way on $\mathcal{O}(U_{N})$ via the following remark:
$c_{ij}(gh)=\sum_{k=1}^{N}c_{ik}(g)c_{kj}(h)=\sum_{k=1}^{N}(c_{ik}\otimes c_{kj})(g,h),\quad g,h\in U_{N},1\leq i,j\leq N.$
The “group law” on $U_{N}^{+}$ is therefore given by the unique *-homomorphism $\Delta:\mathcal{O}(U_{N}^{+})\to\mathcal{O}(U_{N}^{+})\otimes\mathcal{O}(U_{N}^{+})$, called coproduct, such that
$\Delta(u_{ij})=\sum_{k=1}^{N}u_{ik}\otimes u_{kj}.$
The existence of $\Delta$ follows from the universal property. Additionally, let us outline the existence of the counit and the antipode, which are crucial objects. The counit is given by the homomorphism $\varepsilon:\mathcal{O}(U_{N}^{+})\to\mathbb{C}$, defined by $\varepsilon(u_{ij})=\delta_{ij}$. The antipode, on the other hand, is the antihomomorphism $S:\mathcal{O}(U_{N}^{+})\to\mathcal{O}(U_{N}^{+})$, defined by $S(u_{ij})=u_{ji}^{*}$. These maps serve as the unit and the inverse map in the noncommutative setting, respectively.
In this context, probability measures can be generalized by identifying them with their integration linear forms. These forms correspond to states, which are unital positive linear forms on $\mathcal{O}(U_{N}^{+})$. Notably, there exists a particular state that serves as the analog of the Haar measure on $U_{N}^{+}$ (see [36]).
@@
The use of representation theory as a powerful tool to investigate the asymptotic behavior of random walks on groups has been established since the seminal works of P. Diaconis and his coauthors (see, for example, [14, Chap 4]). To describe the representations of $U_{N}^{+}$, we will make use of the free orthogonal quantum group $O_{N}^{+}$, given by the *-algebra that is the quotient $\mathcal{O}(O_{N}^{+})\simeq\mathcal{O}(U_{N}^{+})/\{u_{ij}-u_{ij}^{*}\}$. We will denote by $o_{ij}$ the image of $u_{ij}$ under the quotient map $\mathcal{O}(U_{N}^{+})\to\mathcal{O}(O_{N}^{+})$. The coproduct factors through the quotient map yielding a compact quantum group structure.
If $z$ denotes the identity function on $\mathbb{T}$, then it was proven in [3, Prop 7] that the map $u_{ij}\mapsto o_{ij}z$ extends to an isomorphism of compact quantum groups between $\mathcal{O}(U_{N}^{+})$ and its image in $\mathcal{O}(O_{N}^{+})*\mathcal{O}(\mathbb{T})$. T. Banica proved in [2] that the characters of the irreducible representations of $O_{N}^{+}$ may be labelled by the nonnegative integers $(\chi_{n})_{n\in\mathbb{N}}$ such that
@@
It then follows from the description of the representation theory of free products [35] that the characters of $U_{N}^{+}$ can be recovered as products of characters of $O_{N}^{+}$ and powers of $z$. A precise description is given in [34, Prop 4.3] which we reproduce here.
@@
$\chi^{\epsilon}_{\mathbf{n}}:=z^{[\epsilon_{1}]_{-}}\chi_{n_{1}}z^{\epsilon_{2}}...z^{\epsilon_{p}}\chi_{n_{p}}z^{[\epsilon_{p+1}]_{+}},$
where $\mathbf{n}=(n_{1},...,n_{p})\in\mathbb{N}^{*p}$, $\epsilon=\epsilon_{1}\in\{\pm 1\}$, $[\epsilon]_{-}=\min(\epsilon,0)$, $[\epsilon]_{+}=\max(\epsilon,0)$ and $\epsilon_{i+1}=(-1)^{n_{i}+1}\epsilon_{i}$. The dimensions can be recovered by applying the counit
$d_{\mathbf{n}}:=\varepsilon\big{(}z^{[\epsilon_{1}]_{-}}\chi_{n_{1}}z^{\epsilon_{2}}...z^{\epsilon_{p}}\chi_{n_{p}}z^{[\epsilon_{p+1}]_{+}}\big{)}=P_{n_{1}}(N)...P_{n_{p}}(N).$
We denote by $\mathcal{O}(U_{N}^{+})_{0}$ the algebra generated by the characters and call it the central algebra of $U_{N}^{+}$. An important feature of this subalgebra is the existence of a conditional expectation (see [1, Chap 9]) $\mathbb{E}:\mathcal{O}(U_{N}^{+})\to\mathcal{O}(U_{N}^{+})_{0}$ that leaves the Haar state invariant, let us recall the definition of such an object.
@@
We call generating functional on $U_{N}^{+}$ an hermitian functional $L:\mathcal{O}(U_{N}^{+})\to\mathbb{C}$ that vanishes on $1$ and is positive on $\ker\varepsilon$. There is a one-to-one correspondence between Lévy processes and generating functionals (see [19, Sec 1.5]) via the formulas
@@
The notion of centrality will become relevant in Section 3 as we introduce the quantum unitary Brownian motion as a central Lévy process. Observe that for any generating functional $L$ on $\mathcal{O}(U_{N}^{+})$, the composition $L\circ\mathbb{E}$ also defines a generating functional on $\mathcal{O}(U_{N}^{+})$ which we call the centralized generating functional. Thus, we may consider centralized generating functional and when doing so have in mind that all their information is contained within the restriction of $L$ to the central algebra $\mathcal{O}(U_{N}^{+})_{0}$.

### v3

@@ 1. Introduction
Given a sequence of compact groups $(G_{N})_{N\in\mathbb{N}}$ each equipped with a Markov process $(X_{t}^{(N)})_{t\geq 0}$, we say that it exhibits cutoff at time $(t_{N})$ if the total variation distance to the Haar probability measure $m_{N}$ drops within a window of order $o(t_{N})$, in other words if
$d_{N}\big(t_{N}(1-\epsilon)\big)\underset{N\to\infty}{\longrightarrow}1d_{N}\big(t_{N}(1+\epsilon)\big)\underset{N\to\infty}{\longrightarrow}0,\quad\epsilon>0,$
where $d_{N}(t)=\sup_{A\subset G}|\mathbb{P}[X_{t}\in A]-m_{N}(A)|$ denotes the total variation distance between the Haar probability measure and the process at time $t$. More precisely, if there exists a continuous function $f:\mathbb{R}\to[0,1]$, decreasing from $1$ to 0 such that
$d_{N}\big(t_{N}+cs_{N}\big)\underset{N\to\infty}{\longrightarrow}f(c),\quad c\in\mathbb{R},$
then, the function $f$ is called a cutoff profile of the process.
@@
The first example of cutoff phenomenon comes from the work of P. Diaconis and M. Shahshahani in the 1980s on random transpositions, using representation theory [DS81]. This discovery sparked significant interest, leading to the identification of numerous examples across various contexts. Notable cases include the dovetail shuffle [Ald83, BD92], Glauber dynamics [LLP10], random walks on random graphs [LS10, BLPS18], as well as the Brownian motion on simple compact Lie groups [Mel14]. In this paper, we focus on compact quantum groups, an object initially introduced by S.L. Woronowicz in [Wor98]. The representation theory of compact groups extends naturally to the quantum setting, allowing for the application of similar techniques. This has been studied extensively, starting with finite quantum groups, as explored by J.P. McCarthy in [Mcc17, Mcc19], and extended to infinite compact quantum groups in works such as [Fre18, Fre19].
@@
While determining a cutoff profile is inherently challenging, there are several processes for which such profiles have been established. For instance, in the case of discrete processes such as random walks on finite groups, the random transposition shuffle [Tey19] exhibits the following asymptotic behavior:
@@
$d_{N}\left(\frac{N}{2}(\ln N+c)\right)\underset{N\to\infty}{\longrightarrow}d_{\mathrm{TV}}\left(\mathrm{Pois}(1+e^{-c}),\mathrm{Pois}(1)\right),\quad c\in\mathbb{R},$
@@
Other examples include the lazy random walk on the hypercube [DGM90], the simple exclusion process on the circle [Lac16], the asymmetric simple exclusion process on a segment [BN22], the Ehrenfest Urn model with multiple urns [NT22], the Gibbs Sampler [NT22], random cycles [NT22] and conjugacy invariant random walks on the symmetric group [OTT25].
Very few examples of cutoffs are known on continuous objects (an example other than the Brownian motion on Lie groups is non-negatively curved diffusions [PS25]), and the only objects on which cutoff profiles could be computed are quantum groups. This was done in [FTW21] where the authors compute the profile for the Brownian motion on the quantum orthogonal group $O_{N}^{+}$ and the quantum permutation group $S_{N}^{+}$, as well as for quantum random transpositions.
@@
While the Brownian motion is the most natural continuous process on simple compact Lie groups, defining a clear analog on most compact quantum groups is not straightforward. The orthogonal and permutation quantum groups provide natural processes that can be interpreted as such (coming from analogies between the central generating functionals on $O_{N}^{+}$ and $S_{N}^{+}$, as proved in [CFK14, Thm 10.2] and [FKS16, Thm 10.10], and the classical case for compact Lie groups in [Lia04]). However, extending this concept to other compact quantum groups presents significant challenges.
@@
Our main result is the identification and analysis of the Brownian motion on the unitary quantum group $U_{N}^{+}$, which is detailed in Section 3 and Sections 5-7 respectively, along with the computation of its right limit profile. Specifically, we establish in Theorem 7.3 that, for appropriate extensions $\widetilde{d}_{N}$ of the distances $d_{N}$ in the quantum setting, we have
@@
$\widetilde{d}_{N}\left(N\ln(\sqrt{2}N)+cN\right)\underset{N\to\infty}{\longrightarrow}d_{\mathrm{TV}}\left(\eta_{c}^{r},\nu_{\mathrm{SC}}\right),\quad c\geq 0,$
@@
$\underset{N\to\infty}{\lim\sup}\,\widetilde{d}_{N}\left(N\ln(\sqrt{2}N)+cN\right)\geq d_{\mathrm{TV}}\left(\eta_{c}^{r},\nu_{\mathrm{SC}}\right),\quad c\in\mathbb{R},$
@@
where $\nu_{\mathrm{SC}}$ denotes the semicircle distribution and $\eta_{c}^{r}$ is a mixing measure from a family of (retranslated) free Meixner distributions (see Subsection 2.2 for definitions).
@@
where $\nu_{c}$ represents a limit measure (or state) for the process at time $t_{N}+cs_{N}$. The main difficulty in the quantum setting arises because the process is not absolutely continuous with respect to the stationary distribution whenever $c<c_{0}$ (for some $c_{0}\in\mathbb{R}\cup\{\infty\}$), depending on the scaling of the sequences). This phenomenon, noted in [Bia08, FTW21], illustrates that processes exhibit an atom in certain regions. In our case, the central algebra of the quantum unitary group $U_{N}^{+}$ is not commutative, unlike $O_{N}^{+}$ and $S_{N}^{+}$. Additionally, in the region where absolute continuity is lost, the singular part does not reduce to a single atom, but rather involves a more complex structure that we were unable to fully identify, which is why we can only establish a lower bound in this region.
Acknowledgment
@@ 2. Preliminaries
In this paper, we will focus on compact quantum groups, which may be unfamiliar to probabilist readers. To address this, we will dedicate a preliminary section to defining and presenting key aspects. Specifically, we will introduce free unitary and orthogonal quantum groups, along with relevant results regarding Lévy processes on these objects. A natural approach to studying the cutoff phenomenon on compact quantum groups is through their representation theory, following classical techniques used for both compact Lie groups and finite groups. Notably, the representation theory of compact quantum Lie groups exhibits structural similarities to that of these classical groups, as it arises from combinatorial constructions involving colored non-crossing partitions (see, for instance, [fre23, Chap 4]). Our exposition is adapted from [FTW21] where definitions were formulated for the quantum orthogonal group $O_{N}^{+}$.
@@ 2.1. Free unitary quantum groups
Free unitary quantum groups, a specific class of compact quantum groups, were introduced by S. Wang in [Wan95]. Quantum groups, more broadly, were originally defined by S.L. Woronowicz [Wor98]. These structures involve C*-algebras, fitting their noncommutative topological nature. However, in this article, we will present an alternative definition that emphasizes the algebraic aspects, making it more accessible to non-expert readers.
@@
The universal property ensures the existence of the coproduct *-homomorphism $\Delta:\mathcal{O}(U_{N}^{+})\to\mathcal{O}(U_{N}^{+})\otimes\mathcal{O}(U_{N}^{+})$ defined by
$\Delta(u_{ij})=\sum_{k=1}^{N}u_{ik}\otimes u_{kj},\quad 1\leq i,j\leq N,$
encoding the “group law" of the compact quantum group. Additionally, let us outline the existence of the counit and the antipode, which are crucial objects. The counit is given by the *-homomorphism $\varepsilon:\mathcal{O}(U_{N}^{+})\to\mathbb{C}$, defined by $\varepsilon(u_{ij})=\delta_{ij}$. The antipode, on the other hand, is the *-antihomomorphism $S:\mathcal{O}(U_{N}^{+})\to\mathcal{O}(U_{N}^{+})$, defined by $S(u_{ij})=u_{ji}^{*}$. These maps serve as the unit and the inverse map in the noncommutative setting, respectively.
In this context, probability measures can be generalized by identifying them with their integration linear forms. These forms correspond to states, which are unital positive linear forms on $\mathcal{O}(U_{N}^{+})$. Notably, there exists a particular state that serves as the analog of the Haar measure on $U_{N}^{+}$ (see [Wor98]).
@@
The use of representation theory as a powerful tool to investigate the asymptotic behavior of random walks on groups has been established since the seminal works of P. Diaconis and his coauthors (see, for example, [Dia88, Chap 4]). To describe the representations of $U_{N}^{+}$, we will make use of the free orthogonal quantum group $O_{N}^{+}$, given by the *-algebra that is the quotient $\mathcal{O}(O_{N}^{+})\simeq\mathcal{O}(U_{N}^{+})/\{u_{ij}-u_{ij}^{*}\}$. We will denote by $o_{ij}$ the image of $u_{ij}$ under the quotient map $\mathcal{O}(U_{N}^{+})\to\mathcal{O}(O_{N}^{+})$. The coproduct factors through the quotient map yielding a compact quantum group structure.
If $z$ denotes the identity function on the unit circle $\mathbb{T}$, then it was proven in [Ban97, Prop 7] that the map $u_{ij}\mapsto o_{ij}z$ extends to an isomorphism of compact quantum groups between $\mathcal{O}(U_{N}^{+})$ and its image in $\mathcal{O}(O_{N}^{+})*\mathcal{O}(\mathbb{T})$. T. Banica proved in [Ban96] that the characters of the irreducible representations of $O_{N}^{+}$ may be labeled by the nonnegative integers $(\chi_{n})_{n\in\mathbb{N}}$ such that
@@
It then follows from the description of the representation theory of free products [Wan95] that the characters of $U_{N}^{+}$ can be recovered as products of characters of $O_{N}^{+}$ and powers of $z$. A precise description is given in [VV13, Prop 4.3] which we reproduce here.
@@
$\chi^{\epsilon}_{\mathbf{n}}:=z^{[\epsilon_{1}]_{-}}\chi_{n_{1}}z^{\epsilon_{2}}\cdots z^{\epsilon_{p}}\chi_{n_{p}}z^{[\epsilon_{p+1}]_{+}},$
where $\mathbf{n}=(n_{1},\cdots,n_{p})\in\mathbb{N}^{*p}$, $\epsilon=\epsilon_{1}\in\{\pm 1\}$, $\epsilon]_{-}=\min(\epsilon,0)$, $\epsilon]_{+}=\max(\epsilon,0)$ and $\epsilon_{i+1}=(-1)^{n_{i}+1}\epsilon_{i}$. The dimensions can be recovered by applying the counit
$d_{\mathbf{n}}:=\varepsilon\big(z^{[\epsilon_{1}]_{-}}\chi_{n_{1}}z^{\epsilon_{2}}\cdots z^{\epsilon_{p}}\chi_{n_{p}}z^{[\epsilon_{p+1}]_{+}}\big)=P_{n_{1}}(N)\cdots P_{n_{p}}(N).$
We denote by $\mathcal{O}(U_{N}^{+})_{0}$ the algebra generated by the characters and call it the central algebra of $U_{N}^{+}$. An important feature of this subalgebra is the existence of a conditional expectation (see [AP, Chap 9]) $\mathbb{E}:\mathcal{O}(U_{N}^{+})\to\mathcal{O}(U_{N}^{+})_{0}$ that leaves the Haar state invariant, let us recall the definition of such an object.
@@
We call generating functional on $U_{N}^{+}$ an hermitian functional $L:\mathcal{O}(U_{N}^{+})\to\mathbb{C}$ that vanishes on $1$ and is positive on $\ker\varepsilon$. There is a one-to-one correspondence between Lévy processes and generating functionals (see [FS16, Sec 1.5]) via the formulas
@@
The notion of centrality will become relevant in Section 3 as we introduce the quantum unitary Brownian motion as a central Lévy process. Observe that for any generating functional $L$ on $\mathcal{O}(U_{N}^{+})$, the composition $L\circ\mathbb{E}$ also defines a generating functional on $\mathcal{O}(U_{N}^{+})$ which we call the centralized generating functional. Thus, we may consider centralized generating functionals and when doing so have in mind that all their information is contained within the restriction of $L$ to the central algebra $\mathcal{O}(U_{N}^{+})_{0}$.